OpenStudy (anonymous):

the Function h(m)=-3/20m+12 Describes the height of a 12-inch candle after it has been burning m Minutes. How many minutes has a 12-inch candle been burning, if it is now 7.5 inches high.

OpenStudy (mathteacher1729):

h(m) should be read as "height, as a function of minutes" or h(m) means "the height of the candle after m minutes have passed" In this case you're looking for the height (which is in inches) so that means h(m) = 7.5 Your goal is to find the number of minutes so... 7.5 = -3/20m+12

OpenStudy (anonymous):

how do you solve that?

OpenStudy (mathteacher1729):

It is a linear equation. Sometimes called a "two step" equation. 7.5 = -3/20m+12 7.5 -12 = -3/20m+12 -12 subtract 12 from both sides -4.5 = -3/20m + 0 -4.5 = -3/20m (-4.5)*20 = (-3/20)*20m multiply both sides by 20 -90 = -3m -90/-3 = -3m/3 divide both sides by 3 30 = m You may say "um, you did THREE steps..." but multiplying by 20 and dividing by 3 can be done in a single step... just multiply by -20/3 and that's it.

OpenStudy (anonymous):

thank you.

OpenStudy (mathteacher1729):

You're welcome. :)